OpenStudy (anonymous):

WILL FAN AND GIVE MEDAL! Multiply [sorry I get confused by these a lot] (–4) ∙ (–8)

OpenStudy (anonymous):

Please help!!!!!!!!!!!!!!!

OpenStudy (shinalcantara):

when a negative number is multiplied by a negative number, the answer is positive

OpenStudy (shinalcantara):

In multiplying numbers, (-)*(-) = + (-)*(+) = - (+)*(-) = - (+)*(+) = +

OpenStudy (anonymous):

thank you so much!

OpenStudy (shinalcantara):

yw